a ladder resists against a frictionless vertical wall,with its upper end 6m above the ground and the lower end 4m below away from the waall.The weight of ladder is 500N and its centre of gravity at 1/3 distance from the lower end.Walls reaction will be in newton 
 

Grade:12

1 Answers

Muthukumar
6 years ago
There is no net force along the ladder. Resolving weight and normal force along the ladder you get N sin \theta = mg cos \theta or N = mg tan\theta, where \theta is the angle made by the ladder with the ground. Substituting the appropriate values, you get N= 750 Newtons.